OpenStudy (anonymous):

the derivative of f is given as 2(x-1)(2x-1)^2(5x-2). f has a relative maximum at a. x=1 only b. x=2/5 only c. x=1/7 only d. x=-1/2 only e. x=1, x=-1/2

OpenStudy (kc_kennylau):

Extrema of f(x) is/are the solution(s) of f'(x)=0

OpenStudy (anonymous):

so do i expand the given derivative? or am i able to solve with the function as is?

OpenStudy (kc_kennylau):

If f''(x) is positive, x is a minimum. If f''(x) is negative, x is a maximum.

OpenStudy (kc_kennylau):

If f''(x) is zero, check your calculation xD

OpenStudy (kc_kennylau):

@laniraegreen just solve the equation f'(x)=0 :) (P.S. f'(x) is given)

OpenStudy (anonymous):

but to solve for an x value, do i need to fully multiply the given f'(x)?

OpenStudy (kc_kennylau):

no you don't, f'(x) is already factorized to facilitate you solving it :)

OpenStudy (kc_kennylau):

If something=0 and "something" is factorized, you set each factor to zero.

OpenStudy (kc_kennylau):

That means x-1=0 or 2x-1=0 or 5x-2=0

OpenStudy (anonymous):

okay! i understand that now.

OpenStudy (kc_kennylau):

remember to plug the solutions to f''(x) to see if it's a minimum or a maximum :)

OpenStudy (anonymous):

so i have my three solutions, but do i differentiate each factor to get f''?

OpenStudy (anonymous):

i know how to factor, i just don't know the next step

OpenStudy (anonymous):

differentiate* haha sorry

OpenStudy (kc_kennylau):

You either: use the product rule; or: expand it then differentiate

OpenStudy (kc_kennylau):

get it?

OpenStudy (anonymous):

i'm just substituting the solutions into f''.....

OpenStudy (kc_kennylau):

I have to go now, so hopefully someone will help you :) @ganeshie8 @Zarkon

OpenStudy (anonymous):

okay, thank you so much for all your help though! :)

OpenStudy (anonymous):

once i get my values, then what? is the negative value the maximum?

ganeshie8 (ganeshie8):

since f'(x) is in factored form, you may try and visualize what intervals f'(x) is increasing/decreasing and figure out if f"(x) will be +ve/-ve/0

ganeshie8 (ganeshie8):

2(x-1)(2x-1)^2(5x-2) degree = 4, use the precalculus concepts. wat can you tell about end-behavior ?

OpenStudy (anonymous):

that both ends will extend into either positive or negative infinity?

ganeshie8 (ganeshie8):

yup ! since the coefficient is +ve, both ends go UP

OpenStudy (anonymous):

so theres one maximum?

ganeshie8 (ganeshie8):